A proton is released in a uniform electric field, and it experiences an electric force of 1.50
Answer #1

F = qE

E = 1.5*10^-14/(1.6*10^-19)

= 93750 N/C

b) South
Answer #2

F = qE

E = 1.5*10^-15/(1.6*10^-19)

= 93750 N/C

or 93.750 kN/C

b) South
Answer #3

F = q*E

==> E = F/q = 1.5*10^-15/1.6*10^-19v = 9375 N/c

direction of E is towards south
Answer #4

F = Eq

E = F/q = 1.5*10^-14 / (1.6*10^-19) = 9.375 * 10^4 N/C


Direction is South because Force direction is South


linear thermal expansion coef brass 19e-6 /K

?L/L = ??T, ? is linear thermal expansion coef
1.7% is 0.017
0.017 = 19*10^-6 * ?T
?T = 894.73
so T = 894.73 +20 = 914.73 C

You can get different numbers, as the coef will vary with different types of brass.